OpenStudy (anonymous):

what is the exact circumference of a circle with the diameter of 25inches

Parth (parthkohli):

Hi. Welcome to OpenStudy! You first go about finding the radius. Can you do that?

Parth (parthkohli):

You divide the diameter by 2.

OpenStudy (anonymous):

circumference of a circle = 2rpi = pi d

Parth (parthkohli):

Then, you use the formula \(\text{Circumference} = 2\pi r\)

Parth (parthkohli):

Or, you could directly compute \(\pi \times d\).

OpenStudy (anonymous):

now subsitute d = 25 inches

Parth (parthkohli):

So, what is \(25 \times 3.14\)? @black12

OpenStudy (anonymous):

I think 25*3.14 will not give the exact answer. It will only give the approximate value. so the exact value is simply 25 pi inches.

Parth (parthkohli):

That is a point too :)